OpenStudy (anonymous):

Which three ordered pairs complete the table to give solutions of the equation y = x – 5? A. (–3, –2), (0, –5), and (3, 2) B. (–3, –2), (0, –5), and (3, –2) C. (–3, –8), (0, –5), and (3, 2) D. (–3, –8), (0, –5), and (3, –2)

OpenStudy (bohotness):

what do you think the answer is?

OpenStudy (anonymous):

I Think its B But it a Way I think its not..

OpenStudy (anonymous):

why do you think its B?

OpenStudy (anonymous):

Because i looked up a Website that Couldve helped me Abd I Found the cordinates And im Not Sure if i Got them In the right Format..

OpenStudy (anonymous):

well the answer is not B, do you want help with understanding this?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

ok so answer some questions for me

OpenStudy (anonymous):

Ok But Im not so good at this Soo..

OpenStudy (anonymous):

thats ok, nobody is good at it until they practice it

OpenStudy (anonymous):

thats the definition of learning

OpenStudy (anonymous):

ok heres my first question

OpenStudy (anonymous):

Yea I Was sick the Day i Was suppose to Practice It and im Tyring to Understand it for today

OpenStudy (anonymous):

if i give you the point (-3, -2) what is the x coordinate and what is the y coordinate of that point?

OpenStudy (anonymous):

Wouldn't It Be..

OpenStudy (anonymous):

Sorry i Crashed

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

ok so whats the answer to my question?

OpenStudy (anonymous):

Wouldve -2 Be the y Cordinate?

OpenStudy (anonymous):

Either that Or im Just Plain COnfused D:

OpenStudy (anonymous):

yes -2 is the y coordinate, what is the x coordinate?

OpenStudy (bohotness):

A -2 = -3 - 5, -2 = -8 is wrong which means A is not the answer B -2 = -3 -5, -2 = -8 is wrong which means B is not the answer C -8 = -3 - 5, -8 = -8 -5 = 0 - 5, -5 = -5 2 = 3 - 5, 2 = -2 is wrong so C is not the answer

OpenStudy (anonymous):

. . .

OpenStudy (anonymous):

@iiForeverLost plz ignore that

OpenStudy (anonymous):

-3 Would be the X Cordinate

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

Yay im Learning

OpenStudy (anonymous):

ok so if i give you (-3, -2) you told me x coordinate is -3 and y coordinate is -2, that is correct, perfect so are you ok if i say x = -3 and y = -2 for that point? using = instead of the word "is"?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

It Makes more sence honestly

OpenStudy (anonymous):

ok perfect, so we have the equation for a line we have y = x - 5 and we want to know if those points are on that line, in order to do that we can replace x with the x coordinate of the point and y with the y coordinate of the point, and then see if that makes a true statement, of that is a true statement then we can plot the point which will be part of our line

OpenStudy (anonymous):

but Wouldnt the Answer Be D Because If X Is -3 And Y Is 2 And x - 5 Is (-5 and 0?

OpenStudy (anonymous):

so lets try plugging the point (-3, -2) into y = x - 5

OpenStudy (anonymous):

im not sure what you mean by that exactly

OpenStudy (anonymous):

lets just go step by step

OpenStudy (anonymous):

Like What i mean is If you connected all the Cordinates they would End up in the order

OpenStudy (anonymous):

umm yes if you connect all the coordinates with a line, they will make a straight line

OpenStudy (anonymous):

for a linear equation

OpenStudy (anonymous):

Exactly And i Went on the COnfusing website again and I Found the cordinates But Im not sure if i did this right..

OpenStudy (anonymous):

ok but forget all that, just focus on what im asking, and we will learn this step by step

OpenStudy (anonymous):

Ok

OpenStudy (bohotness):

-.- plug all your points in love if you have points

OpenStudy (anonymous):

Ok Ive done that and it got Me D..

OpenStudy (bohotness):

:)

OpenStudy (anonymous):

ok good for u, correct

OpenStudy (anonymous):

Is D Right?

OpenStudy (anonymous):

Who Wants the medal?

OpenStudy (anonymous):

yes D is right, i hope you understand why its right, i dont care at all about medals, so plz give it to someone else

OpenStudy (bohotness):

yea

OpenStudy (anonymous):

if you arent sure why D is right and want to to know why, then msg me

OpenStudy (anonymous):

Ok Thank you, Both of You